OpenStudy (anonymous):

find the one sided limit f(x) = square root of 1-cos2x /x ,as x approaches 0

OpenStudy (anonymous):

is it \[\sqrt{1-\cos^2(x)}\] or \[\sqrt{1-\cos(2x)}\]

OpenStudy (anonymous):

it is \[\frac{ \sqrt{1- \cos 2x} }{ x }\]

OpenStudy (anonymous):

does it help to know that \[1-\cos(2x)=2\sin^2(x)\]?

OpenStudy (anonymous):

it is 2x not \[x^{2}\]

OpenStudy (anonymous):

some trig identity you forget as soon as you stop looking at trig identities

OpenStudy (anonymous):

how does 1−cos(2x) becomes 2sin^2(x)

OpenStudy (anonymous):

i was trying to prove that but could not .

OpenStudy (anonymous):

the one that starts \[\cos(2x)=1-\sin^2(x)\]

OpenStudy (anonymous):

i guess it depends on how far back you want to go, but that is one of the "double angle" formulas for cosine

OpenStudy (anonymous):

but can i use cos^2x + sin^2x=1 identity to derive it?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

no it comes from angle sum identity \[\cos(x+x) = \cos^2 x - \sin^2 x\] then you can use pythagorean identity cos^2 = 1-sin^2

OpenStudy (anonymous):

unless it was \[\sqrt{1-\cos^2(x)}\] in which case the answer is different i thought you said it was \[\sqrt{1-\cos(2x)}\]

OpenStudy (anonymous):

like i said, depends on how far back you want to do

OpenStudy (anonymous):

Thank you i have understood now.

OpenStudy (anonymous):

yw